# Energetics-based model for a diffusiophoretic motion of a deformable droplet

### v1

@@ IV Reduction to ODE model including only 2-mode deformation
It is notable that $F^{(x)}_{1}>0$, $F^{(x)}_{3}<0$, $\tilde{F}^{(x)}_{1}=\tilde{F}^{(y)}_{1}<0$, $G_{2}^{(2a)}<0$, $\tilde{G}_{0}^{(2a)}=\tilde{G}_{0}^{(2b)}>0$ for all $R>0$. The other coefficients $\tilde{F}_{3}^{(x)}$, $\tilde{F}_{3}^{(y)}$, $\tilde{G}_{2}^{(2a)}$, and $\tilde{G}_{2}^{(2b)}$ change their signs depending on $R$.
@@ V Analysis of ODE model
$\eta_{t}\frac{d\bm{v}_{c}}{dt}=$ $(-\eta_{t}+f_{1})\bm{v}_{c}-\tilde{f}_{1}S\bm{v}_{c}-f_{3}\left|\bm{v}_{c}\right|^{2}\bm{v}_{c}$
@@
Here, $I$ is an identity tensor and $S$ is a traceless tensor that represents the second-order deformation, which is connected with $a_{2}$ and $b_{2}$ [50, 52] as
@@
Considering the terms up to the third order of small parameters $\bm{v}_{c}$ and $S$, we obtain
@@
Therefore, we finally obtain the ordinary differential equation (ODE) model as
@@
Based on the ODE model, we performed the analyses of the solutions and their stability. The details are shown in Appendix D.
We adopted the coefficient values that corresponded to the numerical simulation ($R=1$) in Section III and calculated the phase diagram shown in Fig. 4. The diagram well reproduces the one obtained from the direct numerical simulation shown in Fig. 2. The slight differences should be due to the effect of the smoothing of the surface and space discretization. The most serious difference is that the bistable region between ID and MD droplets exists in the phase diagram in Fig. 4, while it does not exist in the one in Fig. 2. We consider this is because we neglect the higher-order terms in the process of reduction to the ODE model.
Figure 4: Phase diagram obtained by the ODE model. The boundary curves are given as $\eta_{t}=f_{1}$, $\kappa_{2}=\tilde{g}_{0}$, Eq. (112) with Eq. (111), and Eq. (114). IC, ID, and MD indicate an immobile circular droplet, an immobile deformed droplet, and a mobile deformed droplet, respectively.
Therefore, our theoretical analyses on the bifurcation structure and numerical simulation results suggest that an IC droplet becomes unstable and a minor-axis directed moving elliptic droplet becomes stable with decreasing $\eta_{t}$ through a subcritical pitchfork bifurcation. In contrast, an IC droplet becomes unstable and an ID droplet becomes stable with decreasing $\kappa_{2}$ through a supercritical pitchfork bifurcation. With a further decrease in $\kappa_{2}$, an ID droplet becomes unstable and an MD droplet becomes stable. This transition is discontinuous but we could not identify the bifurcation structure since our reduced ODE model does not properly reflect the transition observed in direct numerical simulation.
VI Discussion
@@
In the present paper, we constructed a model for a deformable self-propelled droplet that exhibits self-diffusiophoresis and performed numerical simulations and theoretical analyses to discuss the relation between motion and deformation using the Fourier expansion for the deformation. The analyses based on the Fourier mode expansion were also reported in previous studies. Ohta and Ohkuma proposed a mathematical model in which the velocity of the self-propelled particle and the amplitude of 2-mode deformation are considered [50, 52]. In their model, it is assumed that deformation decays if the particle does not move. In contrast, we derived the coefficients appearing in the Ohta-Ohkuma model by the reduction from a model including a concentration profile. The obtained results suggest both spontaneous translational motion and spontaneous deformation are possible. Especially, they did not discuss an immobile deformed (ID) droplet since they only considered the case with the decay of deformation. Here, we propose that the ID droplet can exist.
@@
As for the pulse dynamics in reaction-diffusion systems, there have been several studies in which the coupling between the pulse shape and motion is concerned. Krischer and Mikhailov reported a traveling pulse in a two-dimensional reaction-diffusion system with volume conservation using numerical simulation. They exhibited a characteristic shape of a traveling solution [69]. Teramoto et al. studied the reaction-diffusion system with one activator and two inhibitors. They analyzed the bifurcation structure based on the center manifold theorem. They clarified the bifurcation structure including the bifurcation from a circular-shaped standing pulse into a deformed traveling pulse solution [70]. Ohta et al. analyzed the reaction-diffusion equation with volume conservation and discussed the dynamics by expanding with Fourier series [71]. They also expanded their study into a three-dimensional system [72]. Our approach is similar to their work in that the shape is described as the Fourier expansion in polar coordinates. Since our reaction-diffusion equation is linear, the analysis is simpler and the analytical approach can be more easily performed.
@@ Acknowledgements.
Appendix A Detailed derivation of $\partial E/\partial\bm{r}_{c}$, $\partial E/\partial a_{k}$, and $\partial E/\partial b_{k}$ in Eqs. (14), (15), and (16)
Here, we show the detailed derivation of Eqs. (14) , (15), and (16). We set $\bm{r}_{c}=x_{c}\bm{e}_{x}+y_{c}\bm{e}_{y}$, and first we calculate $\partial E_{s}/\partial x_{c}$ as
@@
$=$ $\lim_{\delta\to 0}\frac{\iint_{\mathbb{R}^{2}\backslash\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}+\delta\bm{e}_{x}))dA-\iint_{\mathbb{R}^{2}\backslash\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))dA}{\delta}$
@@
In the same manner, we obtain
@@
$=\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))\bm{e}_{x}\cdot\bm{e}_{n}(\bm{r})d\ell,$ (74)
and in the same manner
@@
In the calculation from the second to third lines, we replace $\bm{r}$ with $\bm{r}+\delta\bm{w}_{k}^{(a)}$, where $\bm{w}_{k}^{(a)}$ is given in Eq. (17). The Jacobian matrix regarding this transform is $1+\mathcal{O}(\epsilon^{2})$. The conditions required for the vector field $\bm{w}_{k}^{(a)}$ are that the divergence is 0, i.e., $\nabla\cdot\bm{w}_{k}^{(a)}=0$ and that it satisfies
@@
The incompressible flow profile in Eq. (17) holds the above conditions.
@@ Appendix B Detailed derivation of the steady-state concentration
We also need to consider the continuity condition at the droplet boundary. The results are already given in our previous papers [58, 59, 60] on a self-propelled solid particle with a deformed shape from a circle. Here, we set $a_{2}=\epsilon\tilde{a}_{2}$ and $b_{2}=\epsilon\tilde{b}_{2}$ to clearly show the dependence of the small parameter $\epsilon$. Then, we obtain
@@ Appendix C Fitting of the coefficient of the cubic term of the deformation
Since our model started from the free energy considering the first order of the deformation magnitude, we cannot analytically derive the coefficient of the cubic term with respect to the deformation magnitude. For the ODE approach, we need to know the value of the coefficient of the cubic term. Therefore, we ran the numerical simulation and obtained the coefficient by fitting the simulation data. Since we only needed to consider the transition from an IC droplet to an ID droplet, we did not include the dynamics of the translational motion but only considered the deformation dynamics coupled with the time evolution of the concentration field. Other parameters were set to be the same as described in the main text. The plot of the deformation magnitude $s$ versus $\kappa_{2}$ is shown with red dots in Fig. 5.
Figure 5: Magnitude $s$ of deformation depending on $\kappa_{2}$ obtained by the numerical simulation only considering the deformation dynamics. The simulation results are shown with red dots. The result of the fitting to a function $\kappa_{2}=p_{0}-p_{2}s^{2}$ is shown with a blue line. Here, $p_{0}$ and $p_{2}$ were obtained to be $p_{0}=0.1184$ and $p_{2}=0.1436$ from the fitting. The green broken line shows the curve using the analytically estimated value for $\tilde{g}_{0}$.
@@
with the least square method. Since we needed to focus on the region with small $\left|s\right|$, we only used the data with $0.001<\left|s\right|<0.3$. By the fitting, we obtained $p_{0}=0.1184$ and $p_{2}=0.1436$. The resulting curve is shown with a blue solid line in Fig. 5. From the theoretical analysis, the constant term corresponding to $p_{0}$ is estimated to be $\tilde{g}_{0}$, which is estimated as $\tilde{g}_{0}\simeq 0.1196$ with the corresponding parameters. The curve using this estimated value for $\tilde{g}_{0}$ is also plotted by a green broken line in Fig. 5. The difference between the estimated value of $\tilde{g}_{0}$ and $p_{0}$ may be due to the smoothing effect. For obtaining the phase diagram in Fig. 4, we adopted the estimated value from the theory for $\tilde{g}_{0}$ and $p_{2}\simeq 0.1436$ for $h_{0}$.
@@ Appendix D Details on analyses of the ODE model
Therefore, the curves $\eta_{t}=f_{1}$, $\kappa_{2}=\tilde{g}_{0}$, Eq. (112) with Eq. (111), and Eq. (114) indicate the boundary of a stable IC droplet region when decreasing $\eta_{t}$, the boundary between stable IC and ID droplet regions, the boundary of a stable MD droplet when increasing $\eta_{t}$ or $\kappa_{2}$, and the boundary of the stable ID droplet when decreasing $\eta_{t}$ or $\kappa_{2}$, respectively, as shown in Fig. 4.

### v2

@@ IV Reduction to ODE model including only 2-mode deformation
It is notable that $F^{(x)}_{1}>0$$F^{(x)}_{3}<0$$\tilde{F}^{(x)}_{1}=\tilde{F}^{(y)}_{1}<0$, $G_{2}^{(2a)}<0$, $\tilde{G}_{0}^{(2a)}=\tilde{G}_{0}^{(2b)}>0$ for all $R>0$. The other coefficients $\tilde{F}_{3}^{(x)}$, $\tilde{F}_{3}^{(y)}$, $\tilde{G}_{2}^{(2a)}$, and $\tilde{G}_{2}^{(2b)}$ change their signs depending on $R$.
@@ V Analysis of ODE model
$m\frac{d\bm{v}_{c}}{dt}=$ $(-\eta_{t}+f_{1})\bm{v}_{c}-\tilde{f}_{1}S\bm{v}_{c}-f_{3}\left|\bm{v}_{c}\right|^{2}\bm{v}_{c}$
@@
Here, $I$ is an identity tensor and $S$ is a traceless tensor that represents the second-order deformation, which is connected with $a_{2}$ and $b_{2}$ [54, 55, 56, 76] as
@@
Considering the terms up to the third order of small parameters $\bm{v}_{c}$ and $S$, we obtain the ordinary differential equation (ODE) model:
@@
Figure 3: Analysis of the ODE model in Eqs. (62) and (63). (a) Phase diagram obtained by the ODE model. The boundary curves are given as $\eta_{t}=f_{1}$, $\kappa_{2}=\tilde{g}_{0}$, Eq. (115) with Eq. (114), and Eq. (117). IC, ID, and MD indicate an immobile circular droplet, an immobile deformed droplet, and a mobile deformed droplet, respectively. IC/MD and ID/MD denote the bistable states of IC and MD droplets and ID and MD droplets, respectively. (b) One-dimensional bifurcation diagrams with constant $\eta_{t}$ or $\kappa_{2}$, which are indicated as arrows at the top and right sides of the plot in panel (a). Speed $v$ and deformation $s$ are plotted. (i,ii) $\kappa_{2}$ is varied at (i) $\eta_{t}=0.085$ and (ii) $\eta_{t}=0.095$. (iii-vi) $\kappa_{2}$ is varied at (iii) $\kappa_{2}=0.160$, (iv) $\kappa_{2}=0.130$, (v) $\kappa_{2}=0.118$, and (vi) $\kappa_{2}=0.105$. The red, green, dark blue, and brown lines represent immobile circular (IC), immobile deformed (ID), mobile deformed (MD) droplets moving in the minor-axis direction, and MD droplets moving in the major-axis direction. The thick and thin lines indicate the stable and unstable solutions, respectively. The characters “P” and “SN” mean the pitchfork and saddle-node bifurcations. The vertical dotted lines are drawn to indicate the transition points and correspondence between $v$ and $s$.
@@
Therefore, we finally obtain the ODE model in the polar coordinates as
@@
Based on the ODE model, we perform the analyses of the solutions and their stability. The details are shown in Appendix D.
We adopt the coefficient values that correspond to the numerical simulation ($R=1$) by the partial differential equation (PDE) model in Section III and the phase diagram and the bifurcation diagrams shown in Fig. 3. These diagrams well reproduce the ones obtained from the direct numerical simulation using the PDE model shown in Fig. 2. The amplitudes of $v$ and $s$ in the bifurcation diagrams also meet between the PDE and ODE models. The quantitative correspondence is closer in the parameter region near the bifurcation point from the ID droplet, i.e., $\eta_{t}=f_{1}$ and $\kappa_{2}=\tilde{g}_{0}$. Far from the bifurcation point, the quantitative correspondence becomes worse, but the qualitative one is maintained.
From the bifurcation diagrams in Fig. 3(b), we can determine the bifurcation types between the different modes. An IC droplet becomes unstable and a minor-axis directed moving elliptic droplet becomes stable with decreasing $\eta_{t}$. This transition is classified into a supercritical drift-pitchfork bifurcation for $\kappa_{2}\gtrsim 0.156$ and a subcritical one for $\kappa_{2}\lesssim 0.156$. It should be noted that this drift-pitchfork bifurcation is symmetric in the two dimension. With a decrease in $\kappa_{2}$, an IC droplet becomes unstable and an ID droplet becomes stable through a supercritical pitchfork bifurcation. It should be noted that this bifurcation is also symmetric in the two dimension. With a further decrease in $\kappa_{2}$, an ID droplet becomes unstable and an MD droplet which moves in the minor-axis direction becomes stable. This transition is classified into a supercritical drift-pitchfork bifurcation for $\eta_{t}\gtrsim 0.0925$ and a subcritical one for $0.0781<\eta_{t}\lesssim 0.0925$. For the MD droplet, only the minor-axis directed motion is stable. For $\kappa_{2}\gtrsim 0.156$, it becomes unstable and an IC droplet becomes stable with an increase in $\eta_{t}$ via a supercritical drift-pitchfork bifurcation. For $0.114\lesssim\kappa_{2}\lesssim 0.156$, the MD droplet becomes unstable via a saddle-node bifurcation. The unstable solution of the MD droplet connects the stable MD droplet and the stable IC and ID droplets. For $\kappa_{2}\lesssim 0.114$, the MD droplet becomes unstable and the ID droplet becomes stable via a supercritical drift-pitchfork bifurcation. This suggests the cusp-like structure among the solutions corresponding to the IC, ID and MD droplets. In addition, the bifurcation analysis reveals that the unstable solution corresponding to the MD droplet moving in the major-axis direction, which connects the solutions for IC and ID droplets, is bifurcated from them via supercritical drift-pitchfork bifurcations. The simulation results using the PDE model can be explained by the results of the bifurcation analysis without any contradiction.
VI Discussion and Conclusion
@@
In the present paper, we constructed a model for a deformable self-propelled droplet that exhibits self-diffusiophoresis and performed numerical simulations and theoretical analyses to discuss the relation between motion and deformation using the Fourier expansion for the deformation. The analyses based on the Fourier mode expansion were also reported in previous studies. Ohta and Ohkuma proposed a mathematical model in which the velocity of the self-propelled particle and the amplitude of 2-mode deformation are considered [54, 56]. They propose the model only by considering the symmetric properties of the system. In contrast, we derived the coefficients appearing in the Ohta-Ohkuma model by the reduction from a model including a concentration profile. The obtained results suggest both spontaneous translational motion and spontaneous deformation are possible. In the paper by Ohta and Ohkuma [54], they limited their analysis to the case of no spontaneous deformation, i.e., $\kappa_{2}+\tilde{g}_{0}<0$. In addition, the combination of the signs of the coupling terms between the velocity and deformation was $\tilde{f}_{1}g_{2}<0$, which is different from our results. As the continuing study of [54], Hiraiwa et al. [76] and Tarama et al. [55] proposed the models with the spontaneous deformation, in which the cubic term of $S$ is included. In their models, they also limited their analysis to the case of the coupling terms as $\tilde{f}_{1}g_{2}<0$. Therefore, the bifurcation structure with the bistable states between IC and MD droplets and between ID and MD droplets is unique to our model. Moreover, we can show that an ID droplet can stably exist in the parameter region derived from the PDE model.
@@
Our problem has parallels with the dynamics of a localized pulse in two-dimensional reaction-diffusion systems. There have been several studies in which the coupling between the pulse shape and motion is concerned. Krischer and Mikhailov reported a traveling pulse in a two-dimensional reaction-diffusion system with volume conservation using numerical simulation. They exhibited a characteristic shape of a traveling solution [77]. Teramoto et al. studied the reaction-diffusion system with one activator and two inhibitors. They analyzed the bifurcation structure based on the center manifold theorem. They clarified the bifurcation structure including the bifurcation from a circular-shaped standing pulse into a deformed traveling pulse solution [78]. Ohta et al. analyzed the reaction-diffusion equation with volume conservation and discussed the dynamics by expanding with Fourier series [79]. They also expanded their study into a three-dimensional system [80]. Our approach is similar to their work in that the shape is described as the Fourier expansion in polar coordinates. Since our reaction-diffusion equation is linear, the analysis is simpler and the analytical approach can be more easily performed.
@@ Acknowledgements.
Appendix A Detailed derivation of $\partial E/\partial\bm{r}_{c}$, $\partial E/\partial a_{k}$, and $\partial E/\partial b_{k}$ in Eqs. (14), (15), and (16) in the main text
Here, we show the detailed derivation of Eqs. (14), (15), and (16) in the main text. We set $\bm{r}_{c}=x_{c}\bm{e}_{x}+y_{c}\bm{e}_{y}$, and first we calculate $\partial E_{s}/\partial x_{c}$ as
@@
$=$ $-\lim_{\delta\to 0}\frac{\iint_{\Omega(\bm{r}_{c}+\delta\bm{e}_{x},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))dA-\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))dA}{\delta}$
@@
Here, we invert the integrated region from the first to the second lines. Then, we replace $\bm{r}$ with $\bm{r}+\delta\bm{e}_{x}$ only for the first term of the numerator and adopt the definition of the derivative. In the same manner, we obtain
@@
$=\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))\bm{e}_{x}\cdot\bm{e}_{n}(\bm{r})d\ell.$ (74)
Here, we use
$\nabla\cdot\left(\gamma(u(\bm{r}))\bm{e}_{x}\right)=\nabla\gamma(u(\bm{r}))\cdot\bm{e}_{x}+\gamma(u(\bm{r}))\nabla\cdot\bm{e}_{x}=\nabla\gamma(u(\bm{r}))\cdot\bm{e}_{x}.$ (75)
In the same manner, we obtain
@@
In the calculation, we invert the integrated region from the first to the second lines. From the second to third lines, we replace $\bm{r}$ with $\bm{r}+\delta\bm{w}_{k}^{(a)}$, where $\bm{w}_{k}^{(a)}$ is given in Eq. (17) in the main text. The Jacobian matrix regarding this transform is $1+\mathcal{O}(\epsilon^{2})$. The conditions required for the vector field $\bm{w}_{k}^{(a)}$ are that the divergence is 0, i.e., $\nabla\cdot\bm{w}_{k}^{(a)}=0$ and that it satisfies
@@
The incompressible flow profile in Eq. (17) in the main text holds the above conditions.
@@
Here, we use
$\nabla\cdot\left(\gamma(u(\bm{r}))\bm{w}^{(a)}_{k}\right)=\nabla\gamma(u(\bm{r}))\cdot\bm{w}^{(a)}_{k}+\gamma(u(\bm{r}))\nabla\cdot\bm{w}^{(a)}_{k}=\nabla\gamma(u(\bm{r}))\cdot\bm{w}^{(a)}_{k}$. (82)
@@
$\nabla\cdot\left(\gamma(u(\bm{r})\bm{e}_{x}\right)=\nabla\gamma(u(\bm{r}))\cdot\bm{e}_{x}+\gamma(u(\bm{r}))\nabla\cdot\bm{e}_{x}=\nabla\gamma(u(\bm{r}))\cdot\bm{e}_{x}.$ (84)
@@ Appendix B Detailed derivation of the steady-state concentration
We also need to consider the continuity condition at the droplet boundary. The results are already given in our previous papers [64, 65, 66, 67] on a self-propelled solid particle with a deformed shape from a circle. Here, we set $a_{2}=\epsilon\tilde{a}_{2}$ and $b_{2}=\epsilon\tilde{b}_{2}$ to clearly show the dependence of the small parameter $\epsilon$. Then, we obtain
@@
These explicit forms are obtained by the perturbation method with respect to small parameters $\epsilon$ and $v$. First, the series expansion in (93) is plugged in Eqs. (86) and (89) and the equations are obtained with every order of $\epsilon$ and $v$, which are successively solved. We should consider the condition in which the concentration field $u$ and its gradient $\nabla u$ are continuous at the periphery of $\Omega(\bm{r}_{c},a_{k},b_{k})$ [64, 65, 66, 67].
@@ Appendix C Fitting of the coefficient of the cubic term of the deformation
Since our model started from the free energy considering the first order of the deformation magnitude, we cannot analytically derive the coefficient of the cubic term with respect to the deformation magnitude. For the ODE approach, we need to know the value of the coefficient of the cubic term. Therefore, we run the numerical simulation and obtain the coefficient by fitting the simulation data. Since we only need to consider the transition from an IC droplet to an ID droplet, we do not include the dynamics of the translational motion but only consider the deformation dynamics coupled with the time evolution of the concentration field. Other parameters are set to be the same as described in the main text. The plot of the deformation magnitude $s$ versus $\kappa_{2}$ is shown with red dots in Fig. 4.
Figure 4: Magnitude $s$ of deformation depending on $\kappa_{2}$ obtained by the numerical simulation only considering the deformation dynamics. The simulation results are shown with red dots. The result of the fitting to a function $\kappa_{2}=p_{0}-p_{2}s^{2}$ is shown with a blue line. Here, $p_{0}$ and $p_{2}$ are obtained to be $p_{0}=0.1184$ and $p_{2}=0.1436$ from the fitting. The green broken line shows the curve using the analytically estimated value for $\tilde{g}_{0}$.
@@
with the least square method. Since we need to focus on the region with small $\left|s\right|$, we only use the data with $0.001<\left|s\right|<0.3$. By the fitting, we obtain $p_{0}=0.1184$ and $p_{2}=0.1436$. The resulting curve is shown with a blue solid line in Fig. 4. From the theoretical analysis, the constant term corresponding to $p_{0}$ is estimated to be $\tilde{g}_{0}$, which is estimated as $\tilde{g}_{0}\simeq 0.1196$ with the corresponding parameters. The curve using this estimated value for $\tilde{g}_{0}$ is also plotted by a green broken line in Fig. 4. The difference between the estimated value of $\tilde{g}_{0}$ and $p_{0}$ may be due to the smoothing effect. For obtaining the phase diagram in Fig. 3, we adopt the estimated value from the theory for $\tilde{g}_{0}$ and $p_{2}\simeq 0.1436$ for $h_{0}$.
@@ Appendix D Details on analyses of the ODE model
Therefore, the curves $\eta_{t}=f_{1}$, $\kappa_{2}=\tilde{g}_{0}$, Eq. (115) with Eq. (114), and Eq. (117) indicate the boundary of a stable IC droplet region when decreasing $\eta_{t}$, the boundary between stable IC and ID droplet regions, the boundary of a stable MD droplet when increasing $\eta_{t}$ or $\kappa_{2}$, and the boundary of the stable ID droplet when decreasing $\eta_{t}$ or $\kappa_{2}$, respectively, as shown in Fig. 3(a).
In order to draw the bifurcation diagram, first we obtain the steady-state solution. The steady-state solutions for the IC and ID droplets are given analytically, and those for the MD droplets are given using the bisection method. For the MD droplets, the solutions for the minor-axis and major-axis directed motion are considered; the solutions moving the other directions should not be the steady state solution. Then, we obtain the linearized equations of Eqs. (62) and (63), and calculate the eigenvalues of the Jacobi matrix of the linearized equation, which gives us the information about the stability and the bifurcation type, which are shown in Fig. 3(b).